The volume of a gas is18.4 ml at 24.0 degrees Celsius. What will the volume be at 12.0 degrees Celsius and 0.889 atm?
1
Expert's answer
2020-05-26T14:21:59-0400

P1*V1/T1 = P2*V2/T2

V2 = T2*V1/T1 = 285K*0.0184L/297K = 0.01766 L